OpenStudy (anonymous):

3x - y - 3z if x = -2, y = 1, and z = -2

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

yea. 3(-2)-1-3(-2)=-6-1+6=-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np